OpenStudy (anonymous):

Find x-value such that f(x)>=78 from the func. f(x)=80-10cos((πx)/12)

OpenStudy (eyust707):

\[80 - 10 \cos({\pi x \over 12}) \ge 78 \]

OpenStudy (eyust707):

Then solve for x.

OpenStudy (anonymous):

what do i do after im left with the cos part?